OpenStudy (anonymous):

in one week, Kean ate a total of 8 ounces of grains at breakfast, fourteen ounces of grains at lunch, and sixteen ounces of grains at dinner. Which of the following shows a simplified ratio of his servings of grain at breakfast to his total weekly servings of grain? four colon fifteen four over nineteen fifteen to nineteen nineteen colon four

OpenStudy (anonymous):

4:19

OpenStudy (anonymous):

8/(8+14+16)= 8/38= 4/19

OpenStudy (anonymous):

^ his answer